Tuesday, 22 December 2009

Closed - DNS Issue

We are currently investigating an issue with DNS  within our network.

Further updates will follow regarding this issue soon.

Sunday, 20 December 2009

Telehouse East connectivity

we are experiencing a problem with connectivity to telehouse east currently.  This is having an effect on services from this location, and our connectivity with netservices.

Further updates will be posted once available.

Tuesday, 8 December 2009

Closed - 20/21cn BT Radius Issue

We have confirmed with BT that there are issues with the BT radius servers.

This is being investigated by BT now.

We are expecting an update very shortly.

Closed - 21cn - bras-red1.pe - Peterborough

There is currently a loss of service with bras-red1.pe - Peterborough

We are currently awaiting an update from BT regarding this Bras.

Monday, 7 December 2009

Description: Work on CP2

Description:  Work on CP2

Date/Time: 08/12/2009 02:00


At 2am work will be carried out on one of our central pipe routers (CP2) to relocate it to our new DSL network core.  This will have no long term effect however users connected to this central will be disconnected.  These should automatically reconnect to another available central pipe router.


Many thanks for your understanding in this matter, and for your continued business.


Closed - 20cn Slow Speed Issues

A small number of users are experiencing speed issues with the 20cn network.

Our engineers are currently investigating this issue and we hope to have an update soon.

Thursday, 3 December 2009

Closed - DNS Issue

We are currently aware of a DNS issue affecting the DNS server,
We hope to have this resolved ASAP.

Wednesday, 2 December 2009

21cn Bras-Red3 Wolverhampton Outage

We believe we have seen a re-occurance of issues with BRAS red3 in wolverhampton, affecting service for 21cn broadband users in and around that region.  BT are currently investigating.

Please check here for futher updates.  186k resellers can check if a user is affected by this outage by using the dolphin advanced diagnostics within their management portal.

Tuesday, 1 December 2009

20cn ADSL Issues

We are currently investigating routing issues on our 20cn ADSL network.
We hope to have an update from our engineers shortly.
21cn Bras-Red3 Wolverhampton Outage

There is currently issues with the Bras-Red3 in Wolverhampton.
This is currently under investigation, and we are awaiting an update.

Once we have received more information we will update this post.

Monday, 9 November 2009

Geographical Broadband Outage - Faraday

We have identified a potential geographical broadband outage in areas connected to the Faraday point of presence. This is affecting 20cn and 21cn lines.

Further information to follow

Tuesday, 27 October 2009

[MSO Slough - Reading]

We have identified a major outage in the Slough and Reading area affecting 20cn and 21cn circuits in the Berkshire area.

We are awaiting more information and will update you once received.

[Co-location Customers Leeds]

Morning Outage 27th Oct 2009 (00:30am to 00:45am).

At 00:30 this morning we experienced two 5 minute outages. These outages were a result of emergency firmware changes at "Poplar" (London Docklands) preformed by NTL who supply one of our Leeds-London link;

We received no notification of this planned work but are taking steps to ensure we receive all future notifications.

Co-location customers may have lost connectivity to their servers for the duration of this outage.

We apologise for any inconvenience caused.

Wednesday, 21 October 2009

20cn blip 20:15 - 20:30

At approx 20:15 there was an issue affecting 20cn lines, the affected lines started re connecting at approximately 20:30.

We are currently speaking to BT regarding an explanation of the connection drop.

We will keep you informed.

BRAS outage

We have identified a potential issue with BRAS red1.ls-bas within the 21st century network.

We believe this may be causing issues in the bradford and west yorkshire area.

We are working with BT to determine what may be causing this and come up with a resolution.


A potential issue has been detected with BRAS-RED1 on the Bristol section of the 21st century network.

We are currently investigating alongside BT on this issue. Potential affected exchanges include:


Friday, 16 October 2009


BT are currently working on an issue affecting 21CN-BRAS-RED1.BS (Bristol)


Monday, 5 October 2009

Notice of IP Address changes

Please be aware that throughout today there will be some IP address changes, this will only affect certain customers and any technical issues should be reported to Support as normal.


Thursday, 24 September 2009

Outage for customers

Start Time: 09:03 - 24/09/09
We have had a brief connectivity issue on one of our central pipes effecting end user connections. Ordinarily these connections would have reconnected immediately, however the BT authentication servers were under extraordinary load at this time and as such the reconnection's were slower than would ordinarily be expected.
This issue was cleared at 09:38 - 24/09/09.
All effected users should now be able to reconnect. In some cases the end user may need to reboot their router to re-initiate the connection.
We apologise for any inconvenience caused.

The 186k Team

Friday, 18 September 2009

Current outage affecting services

There is currently a BT Wholesale service outage affecting Slough and the surrounding areas. We are awaiting an update from BT at this time.

The 186k Team